### v1

# Quantifying Interpretability in CLIP Models with Concept Consistency

Abstract
CLIP is one of the most popular foundational models and is heavily used for many vision-language tasks. However, little is known about the inner workings of CLIP. While recent work has proposed decomposition-based interpretability methods for identifying textual descriptions of attention heads in CLIP, the implications of conceptual consistency in these text labels on interpretability and model performance has not been explored. To bridge this gap, we study the conceptual consistency of text descriptions for attention heads in CLIP-like models. We conduct extensive experiments on six different models from OpenAI and OpenCLIP which vary by size, type of pre-training data and patch size. We propose Concept Consistency Score (CCS), a novel interpretability metric that measures how consistently individual attention heads in CLIP models align with specific concepts. To assign concept labels to heads, we use in-context learning with ChatGPT, guided by a few manually-curated examples, and validate these labels using an LLM-as-a-judge approach. Our soft-pruning experiments reveal that high CCS heads are critical for preserving model performance, as pruning them leads to a significantly larger performance drop than pruning random or low CCS heads. Notably, we find that high CCS heads capture essential concepts and play a key role in out-of-domain detection, concept-specific reasoning, and video-language understanding. These results position CCS as a powerful interpretability metric for analyzing CLIP-like models.
1 Introduction
Large-scale vision-language (VL) models such as CLIP  [30] have significantly advanced state-of-the-art performance in vision tasks in recent years. Consequently, CLIP has been extensively used as a foundational model for downstream tasks such as video retrieval, image generation, and segmentation  [26, 25, 4, 13, 22]. This has enabled the construction of compositional models combining CLIP with other foundation models, thereby increasing the functionality of CLIP while also adding complexity to the overall model structure. Furthermore, tools and methods for interpreting the functioning of such models are becoming increasingly important in order to enable their adoption in domains where an understanding of model predictions is required. Therefore, given its widespread use as a foundation model for VL tasks, it is imperative to understand the inner workings of CLIP models.
Recently, Gandelsman et al. [15] proposed an interpretability approach for CLIP-like models which decomposes image representations into a sum over contributions from individual attention heads, layers, and image patches. This enables the identification of text sequences which characterize the semantic role played by different attention heads of CLIP. While this provides valuable insights into the inner workings of CLIP, a critical question remains regarding the conceptual consistency of text labels automatically identified for various attention heads of the model. This question has important implications for the interpretability, as lower levels of consistency in the concepts identified for any individual attention head could make its functional role in the model more difficult to understand.
To address this important question, we introduce a methodology for systematically measuring the conceptual consistency of text descriptions extracted for describing the semantic roles of attention heads in CLIP-like models. We conduct our study using six different models: ViT-B-16, ViT-B-32, and ViT-L-14 from OpenAI, as well as ViT-B-16, ViT-B-32, and ViT-L-14 from OpenCLIP [21]. First, we identify interpretable structures within the individual heads of the last four layers of the model using a set of text descriptions. To accomplish this, we employ the TextSpan algorithm  [15], which helps us find the most appropriate text descriptions for each head. After identifying these text descriptions, we assign labels to each head representing the common property shared by the descriptions. This labeling process is carried out using in-context learning with ChatGPT. We begin by manually labeling five pairs of text descriptions and their corresponding property labels, which serve as examples. These examples are then used to prompt ChatGPT to assign labels for the remaining heads.
Leveraging the resulting text descriptions of attention heads, we introduce the Concept Consistency Score (CCS), a new interpretability metric that quantifies how strongly individual attention heads in CLIP models align with specific concepts. Using GPT-4o as an automatic judge, we compute CCS for each head and classify them into high, moderate, and low categories based on defined thresholds. A key contribution of our work is our targeted soft-pruning experiments which show that heads with high CCS are essential for maintaining model performance; pruning these heads causes a significantly larger performance drop compared to pruning random or low CCS heads. We also show that high CCS heads are not only crucial for general vision-language tasks but are especially important for out-of-domain detection and concept-specific reasoning. Additionally, our experiments in video retrieval highlight that high CCS heads are equally vital for temporal and cross-modal understanding, thereby underscoring the broad relevance of CCS in analyzing and interpreting CLIP-like models.
Figure 1: Figure shows the steps of computing Concept Consistency Score for each head.
High CCS $(CCS=5)$ Moderate CCS $(CCS=3)$ Low CCS $(CCS\leq 1)$
L23.H11 (“People”) L23.H0 (“Material”) L21.H6 (“Professions”)
Playful siblings Intrica wood carvingte Photo taken in the Italian pizzerias
A photo of a young person Nighttime illumination thrilling motorsport race
Image with three people Image with woven fabric design Urban street fashion
A photo of a woman Image with shattered glass reflections An image of a Animal Trainer
A photo of a man A photo of food A leg
L22.H10 (“Animals”) L11.H0 (“Locations”) L10.H6 (“Body parts”)
Image showing prairie grouse Photo taken in Monument Valley A leg
Image with a donkey Majestic animal colorful procession
Image with a penguin An image of Andorra Contemplative monochrome portrait
Image with leopard print patterns An image of Fiji Graceful wings in motion
detailed reptile close-up Image showing prairie grouse Inviting reading nook
L23.H5 (“Nature”) L11.H11 (“Letters”) L9.H2 (“Textures”)
Intertwined tree branches A photo with the letter J Photo of a furry animal
Flowing water bodies A photo with the letter K Closeup of textured synthetic fabric
A meadow A swirling eddy Eclectic street scenes
A smoky plume A photo with the letter C Serene beach sunset
Blossoming springtime blooms awe-inspiring sky Minimalist white backdrop
Table 1: Examples of high, moderate and low CCS heads.
2 Related Work
Early research on interpretability primarily concentrated on convolutional neural networks (CNNs) due to their intricate and opaque decision-making processes [37, 32, 33, 14, 18]. More recently, the interpretability of Vision Transformers (ViT) has garnered significant attention as these models, unlike CNNs, rely on self-attention mechanisms rather than convolutions. Researchers have focused on task-specific analyses in areas such as image classification, captioning, and object detection to understand how ViTs process and interpret visual information [10, 12, 28, 36, 9, 11]. One of the key metrics used to measure interpretability in ViTs is the attention mechanism itself, which provides insights into how the model distributes focus across different parts of an image when making decisions [8, 6]. This has led to the development of techniques that leverage attention maps to explain ViT predictions. Early work on multimodal interpretability, which involves models that handle both visual and textual inputs, probed tasks such as how different modalities influence model performance [5, 27] and how visual semantics are represented within the model [17, 24]. Aflalo et al. [1] explored interpretability methods for vision-language transformers, examining how these models combine visual and textual information to make joint decisions. Similarly, Stan et al. [34] proposed new approaches for interpreting vision-language models, focusing on the interactions between modalities and how these influence model predictions. Our work builds upon and leverages the methods introduced by Gandelsman et al. [15, 16] to interpret attention heads, neurons, and layers in vision-language models, providing deeper insights into their decision-making processes.
Model Model size Patch size Pre-training data Kappa Spearman’s correlation ($\rho$) Kendall ($\uptau$)
CLIP B 32 OpenAI-400M 0.757 0.781 0.781
CLIP B 16 OpenCLIP-LAION2B 0.676 0.678 0.678
CLIP L 14 OpenAI-400M 0.758 0.758 0.758
Table 2: Results between human judgment and LLM judgment on CCS labelling.
Model Model size Patch size Pre-training data CCS@0 CCS@1 CCS@2 CCS@3 CCS@4 CCS@5
CLIP B 32 OpenAI-400M 0.021 0.062 0.167 0.271 0.312 0.167
CLIP B 32 OpenCLIP-datacomp 0.104 0.062 0.208 0.189 0.208 0.229
CLIP B 16 OpenAI-400M 0.021 0.062 0.125 0.292 0.292 0.208
CLIP B 16 OpenCLIP-LAION2B 0.062 0.062 0.105 0.25 0.25 0.271
CLIP L 14 OpenAI-400M 0.062 0.109 0.172 0.204 0.203 0.25
CLIP L 14 OpenCLIP-LAION2B 0.016 0.031 0.109 0.219 0.297 0.328
Table 3: Concept Consistency Score (CCS) for CLIP models.
3 Methodology
In this section, we outline the methodology used in our analysis, beginning with an explanation of the TextSpan algorithm. We then describe how we extend this algorithm to apply it across all attention heads in multiple CLIP models using in-context learning.
The TextSpan algorithm  [15] is designed to decompose individual attention heads by associating them with corresponding text descriptions. It requires an initial set of text descriptions broadly capturing the concepts in an image. The algorithm starts by generating two matrices: $C$, which contains the outputs for a specific head (denoted as $(l,h)$), and $R$, which includes representations of candidate text descriptions projected onto the span of $C$. In each iteration, TextSpan calculates the dot product between each row of $R$ and the outputs in $C$ to identify the row with the highest variance, known as the "first principal component." Once identified, this component is projected away from all rows, and the process is repeated to find additional components. The projection step ensures that each new component adds variance that is orthogonal to the earlier components, thereby isolating distinct aspects of the text descriptions relevant to each head.
TextSpan is effective at identifying text descriptions that are most similar to a given head. Note that, TextSpan associates only the text descriptions to a head and doesn’t label the common concept. In the next section, we explain in detail our approach to label the concept learnt by individual heads in CLIP models.
4 Quantifying interpretability in CLIP models
4.1 Concept Consistency Score (CCS)
We introduce the Concept Consistency Score (CCS) as a systematic metric for analyzing the concepts (properties) learned by transformer layers and attention heads in CLIP-like models. This score quantifies the alignment between the textual representations produced by a given head and an assigned concept label. Figure  1 illustrates our approach, with the following sections detailing each step in computing CCS.
4.1.1 Extracting Text Representations
From each layer and attention head of the CLIP model, we obtain a set of five textual outputs, denoted as $\{T_{1},T_{2},T_{3},T_{4},T_{5}\}$, referred to as TextSpans. These outputs serve as a textual approximation of the concepts encoded by the head.
4.1.2 Assigning Concept Labels
Using in-context learning with ChatGPT, we analyze the set of five TextSpan outputs and infer a concept label $C_{h}$ that best represents the dominant concept captured by the attention head $h$. This ensures that the label is data-driven and reflects the most salient pattern learned by the head.
4.1.3 Evaluating Concept Consistency
To assess the consistency of a head with respect to its assigned concept label, we employ a state-of-the-art vision-language model, GPT-4o, as an external evaluator. For each TextSpan $T_{i}$ associated with head $h$, GPT-4o determines whether it aligns with the assigned concept $C_{h}$. The Concept Consistency Score (CCS) for head $h$ is then computed as:
$\text{CCS}(h)=\sum_{i=1}^{5}\mathbb{1}\left[T_{i}\text{ aligns with }C_{h}\right]$
where $\mathbb{1}[\cdot]$ is an indicator function that returns 1 if GPT-4o judges $T_{i}$ to be consistent with $C_{h}$, and 0 otherwise.
We define $CCS@K$ as the fraction of attention heads in a CLIP model that have a Concept Consistency Score (CCS) of $K$. This metric provides a global measure of how many heads strongly encode interpretable concepts. A higher $CCS@K$ value indicates that a greater proportion of heads exhibit strong alignment with a single semantic property. Mathematically, $CCS@K$ is defined as:
$CCS@K=\frac{1}{H}\sum_{h=1}^{H}\mathbb{1}\left[\text{CCS}(h)=K\right]$
where $H$ is the total number of attention heads in the model, $\text{CCS}(h)$ is the Concept Consistency Score of head $h$, $\mathbb{1}[\cdot]$ is an indicator function that returns 1 if $\text{CCS}(h)=K$, and 0 otherwise. This metric helps assess the overall interpretability of the model by quantifying the proportion of heads that consistently capture well-defined concepts. Table  1 shows the examples of heads with different CCS scores.
By computing CCS across all attention heads and layers in a CLIP model, we can systematically interpret how different heads specialize in capturing semantic concepts. This provides crucial insights into the internal representations learned by CLIP, revealing whether specific concepts (e.g., object categories, attributes, locations) are localized within certain heads or distributed across multiple layers. This approach advances model interpretability by offering a quantifiable and automated means to analyze the emergent concepts within CLIP’s attention mechanisms, bridging the gap between learned representations and human-understandable concepts.
4.2 Evaluating LLM Judgment Alignment with Human Annotations
In the previous section, we introduced the Concept Consistency Score (CCS), which is computed using GPT-4o as an external evaluator. However, this reliance on a large language model (LLM) for judgment raises a critical question: Are LLM evaluations reliable and aligned with human assessments? To address this, we conduct a human evaluation study to systematically compare LLM judgments with human annotations.
To assess the reliability of LLM-based evaluations, we collect 50 TextSpan descriptions from three different models, along with their corresponding concept labels assigned during CCS computation. One of the authors of this work then evaluated whether each TextSpan is semantically related to its assigned concept label. The collected human ratings serve as the ground truth against which LLM judgments are compared. To measure the alignment between human and LLM evaluations, we compute the following statistical agreement metrics: Cohen’s Kappa, Spearman’s Rank Correlation ($\rho$) and Kendall’s Tau ($\uptau$).
Table 2 presents the computed agreement scores. Notably, the Cohen’s Kappa values exceed 0.65, indicating substantial agreement between human and LLM evaluations. Similarly, the Spearman’s correlation and Kendall’s Tau values are consistently above 0.65, demonstrating a positive correlation between LLM and human judgments. These results provide strong empirical support for using LLMs as evaluators in concept consistency analysis. The high alignment between human and LLM judgments indicates that LLMs can reliably assess semantic coherence, reducing the need for costly and time-consuming human annotations. In the next section, we introduce tasks and the corresponding datasets used in our experiments.
4.3 Experimental Setting
4.3.1 Tasks
We primarly focus our evaluation on image classification, out-of-domain classification and video retrieval tasks. Below, we list the evaluation datasets for each of the tasks.
Image classification: CIFAR-10  [23], CIFAR-100  [23], Food-101  [3], Country-211  [30] and Oxford-pets  [29].
Out-of-domain classification: Imagenet-A  [20] and Imagenet-R  [19].
Video retrieval: MSRVTT  [35], MSVD  [7], DiDeMo  [2].
4.3.2 Models
For experiments we use the following six foundational image-text models: ViT-B-32, ViT-B-16 and ViT-L-14 pretrained from OpenAI-400M  [30] and LAION2B  [31]. Next, we discuss in detail the results from the experiments.
Model CIFAR-10 CIFAR-100 FOOD-101
Original High CCS Low CCS Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 75.68 71.31 73.61 65.08 56.07 62.39 84.01 73.42 82.12
ViT-B-32-datacomp 72.07 70.50 70.43 54.95 53.14 53.72 41.66 38.13 40.77
ViT-B-16-OpenAI 78.10 63.93 76.44 68.22 51.70 65.38 88.73 76.35 87.36
ViT-B-16-LAION 82.82 78.91 75.38 76.92 65.55 72.51 86.63 67.54 81.4
ViT-L-14-OpenAI 86.94 86.29 85.97 78.28 75.66 77.55 93.07 90.75 92.79
ViT-L-14-LAION 88.29 86.48 88.19 83.37 80.07 83.25 91.02 86.45 90.35
Table 4: Accuracy comparison of various CLIP models on CIFAR-10, CIFAR-100 and FOOD-101 datasets. The values represent original accuracy, performance after pruning high-CCS heads, and performance after pruning low-CCS heads.
Model Country-211 Oxford-pets
Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 17.16 11.46 16.3 50.07 46.66 48.96
ViT-B-32-datacomp 4.43 4.37 4.37 26.48 25.98 25.33
ViT-B-16-OpenAI 22.81 10.72 21.79 52.72 49.12 51.89
ViT-B-16-LAION 20.45 7.49 16.87 65.79 48.48 49.81
ViT-L-14-OpenAI 31.91 23.21 30.63 61.79 62.04 62.08
ViT-L-14-LAION 26.41 16.38 25.66 54.1 56.12 57.16
Table 5: Accuracy comparison of various CLIP models on Country-211 and Oxford-pets datasets. The values represent original accuracy, performance after pruning high-CCS heads, and performance after pruning low-CCS heads.
(a) CIFAR-10
(b) CIFAR-100
(c) Food-101
Figure 2: Zero-shot performance comparison for CIFAR-10, CIFAR-100, and Food-101 datasets under different pruning strategies. For random pruning, results are averaged across three runs.
Model Imagenet-A Imagenet-R
Original High CCS Low CCS Original High CCS Low CCS
ViT-B-32-OpenAI 31.49 20.24 28.72 69.09 54.47 64.45
ViT-B-32-datacomp 4.96 4.59 4.65 34.06 31.6 32.47
ViT-B-16-OpenAI 49.85 25.49 47.27 77.37 55.52 74.84
ViT-B-16-LAION 37.97 25.27 27.44 80.56 66.32 71.73
ViT-L-14-OpenAI 70.4 68.15 69.2 87.87 86.56 86.97
ViT-L-14-LAION 53.8 42.44 52.93 87.12 82.22 86.94
Table 6: Accuracy comparison of various CLIP models on ImageNet-A and ImageNet-R. The values represent original accuracy, performance after pruning high-CCS heads, and performance after pruning low-CCS heads.
Figure 3: Zero-shot results on Country-211 (location) dataset.
Figure 4: Zero-shot results on CIFAR-10 (Objects) dataset.
5 Results and Discussion
5.1 Concept Consistency Scores (CCS) for CLIP models.
We measure $CCS@K$ for all values of $K$ i.e $K\in[0,5]$. Table 3 presents the Concept Consistency Score (CCS) distribution across various CLIP models, categorized by architecture size, patch size, and pre-training data. Several noteworthy trends emerge from this analysis. First, models pre-trained on larger and more diverse datasets (e.g., OpenCLIP-LAION2B) tend to exhibit a higher proportion of heads with CCS@5, indicating that a greater number of transformer heads are aligned with semantically meaningful concepts. For instance, the ViT-L-14 model trained on LAION2B shows the highest CCS@5 score of 0.328, suggesting that approximately 32.8% of heads are consistently associated with a single concept, reflecting strong concept alignment in these models.
Second, smaller models such as ViT-B-32 trained on OpenAI-400M demonstrate a significantly lower CCS@5 score (0.167) and a higher proportion of heads with lower CCS values (e.g., CCS@0 = 0.021), indicating weaker alignment of heads to consistent concepts. This observation implies that larger models with richer pre-training data are better at learning concept-specific representations, a key requirement for robust and interpretable multimodal reasoning.
Interestingly, when comparing models with the same architecture but different pre-training corpora, such as ViT-B-32 (OpenAI-400M vs. OpenCLIP-datacomp), we observe a higher CCS@5 score for datacomp (0.229) than OpenAI-400M (0.167), suggesting that dataset composition significantly affects the emergence of interpretable heads.
Moreover, progressive increases in CCS from CCS@0 to CCS@5 show how concept alignment varies within each model. For instance, while ViT-L-14 (OpenCLIP-LAION2B) has a low CCS@0 of 0.016, it steadily increases to a high CCS@5 of 0.328, suggesting that although a few heads are poorly aligned, a substantial fraction are highly consistent in capturing specific concepts.
In summary, these results demonstrate that the CCS metric effectively captures differences in conceptual alignment across models of varying size and pre-training datasets. Models with larger capacities and richer pre-training datasets tend to exhibit higher concept consistency, offering better interpretability and potentially stronger generalization abilities. This analysis underscores the value of CCS as a diagnostic tool for evaluating and comparing the internal conceptual representations learned by CLIP-like models.
5.2 Interpretable CLIP Models: The Role of CCS.
In the previous section, we analyzed the distribution of CCS scores across various CLIP models. This naturally raises an important question: How does CCS provide deeper insights into the functional role of individual attention heads in influencing downstream tasks? To address this, we conduct a soft-pruning analysis targeting attention heads with extreme CCS scores — specifically, heads with high CCS scores (i.e., CCS = 5) and low CCS scores (CCS $\leq$ 1). Soft-pruning is implemented by zeroing out the attention weights of selected heads, thereby nullifying their contribution to the model’s predictions without altering the overall model architecture.
Tables  4 and  5 report zero-shot classification performance on CIFAR-10, CIFAR-100, FOOD-101, Country-211, and Oxford-Pets, under three conditions: no pruning (original), pruning of high-CCS heads, and pruning of low-CCS heads. We observe that pruning heads with high CCS scores consistently leads to a significant degradation in performance across all datasets, whereas pruning low CCS heads has a much smaller impact. This clear difference in sensitivity demonstrates that CCS effectively identifies heads that capture critical, concept-aligned information essential for model predictions. Thus, CCS serves as a reliable indicator of the most functionally important heads in CLIP models, offering a concrete step toward interpreting the internal mechanisms of these large vision-language models.
Interestingly, we also observe variations in sensitivity across model architectures. The ViT-B-16 models experience the most substantial performance drops when high CCS heads are pruned, suggesting a higher dependency on a smaller set of concept-specialized heads. In contrast, ViT-L-14 models exhibit the least drop, indicating a more distributed representation. Among smaller models, OpenAI-trained models show a larger performance degradation compared to OpenCLIP counterparts when high CCS heads are removed, whereas for larger models like ViT-L-14, OpenCLIP models exhibit a slightly higher drop than OpenAI models. These trends further highlight how CCS can reveal architectural and training-specific differences in how CLIP models encode and utilize concepts for downstream tasks.
5.3 High CCS vs random heads pruning
To rigorously evaluate the effectiveness of the Concept Consistency Score (CCS) as a measure of interpretability in CLIP models, we compare the impact of pruning heads with high CCS scores against pruning an equal number of randomly selected heads. While earlier results demonstrated that pruning high CCS heads significantly degrades model performance, a critical question remains: Are these high-CCS heads genuinely more important than other heads?
To investigate this, we conduct controlled experiments where we randomly prune the same number of attention heads and analyze the corresponding performance drop. The results of these experiments are presented in Figure 2. As shown in the figure, pruning high CCS heads consistently leads to a substantially larger decrease in zero-shot performance compared to random pruning, across different datasets and model variants. These findings empirically validate that CCS effectively identifies heads that are essential for the model’s decision-making process, thereby offering a principled mechanism for interpreting the internal workings of CLIP models. Unlike random pruning, which affects heads without regard to their learned properties, CCS-guided pruning systematically targets heads that encode critical concepts, revealing their functional role in model predictions.
Moreover, we observe that larger CLIP models exhibit smaller differences between high CCS and random pruning impacts compared to smaller models, suggesting that larger architectures may possess more redundancy or distributed representations, making them more resilient to head pruning. Overall, these results establish CCS as a reliable and interpretable metric for identifying concept-relevant heads, contributing to a deeper understanding of how CLIP models organize and utilize conceptual knowledge across their layers.
5.4 High CCS heads are crucial for out-of-domain (OOD) detection
While our earlier experiments primarily focused on in-domain datasets such as CIFAR-10 and CIFAR-100 to validate the Concept Consistency Score (CCS), understanding model behavior under out-of-domain (OOD) conditions is a critical step toward evaluating models’ robustness and spurious correlations. Table  6 demonstrates the results on ImageNet-A and ImageNet-R datasets respectively. From the table, we observe that pruning heads with high CCS scores leads to a substantial degradation in model performance, underscoring the critical role these heads play in the model’s decision-making process. Notably, the ViT-B-16-OpenAI model exhibits the most pronounced drop in performance upon pruning high CCS heads, suggesting that this model relies heavily on a smaller set of concept-specific heads for robust feature representation consistent with the observations previously. These results demonstrate that CCS is a powerful metric for identifying attention heads that encode essential, generalizable concepts in CLIP models while avoiding spurious correlations.
(a) MSRVTT
(b) MSVD
(c) DIDEMO
Figure 5: Zero-shot performance comparison of unpruned (original) model, pruning high CSS, low CSS and random heads on video retrieval task.
5.5 Impact of High CCS heads on corresponding tasks.
To further investigate the functional role of high CCS heads, we perform concept-specific pruning experiments. Specifically, we prune heads with high CCS scores for a target concept (e.g., locations) and evaluate the model on tasks aligned with that concept (e.g., location classification). To contrast this, we also prune heads associated with unrelated concepts (e.g., animals) and compare the impact on task performance. The motivation behind this analysis is to validate whether CCS-identified heads are not only aligned with specific concepts but also crucial for performing tasks involving those concepts. Figure  3 and  4 shows the zero-shot results on Country-211 and CIFAR10-objects datasets. These datasets primarily test varied classes of countries and objects respectively.
5.5.1 High CCS heads are crucial for concept-aligned tasks.
Pruning heads with high CCS scores for a target concept (e.g., locations) leads to a significant drop in task performance when the task is aligned with that concept. For example, in ViT-B-16 (openai), location classification accuracy drops from 22.81% to 14.09% when location heads are pruned. This clearly shows that CCS accurately identifies heads encoding concept-relevant information, making them essential for corresponding tasks.
5.5.2 Minimal Impact of Unrelated Head Pruning.
When pruning heads associated with concepts unrelated to the task (e.g., pruning animal heads for location classification), the model’s performance is largely unaffected, demonstrating the concept specificity of high CCS heads.
5.5.3 Objects heads are the most critical for general classification.
In general classification tasks, pruning object-related heads consistently reduces accuracy more than pruning location or color heads. For example, In ViT-B-32 (openai), accuracy drops from 87.6% to 86.02% with object head pruning, while location and color head pruning show smaller drops (87.02% and 87.22%, respectively). This indicates that object heads are more critical to overall recognition tasks, aligning with the importance of object features in vision models.
5.5.4 Larger models show robustness to pruning.
Larger models like ViT-L-14 exhibit smaller performance drops when pruning concept-specific heads, suggesting more distributed and redundant representations. For example, In ViT-L-14 (openai), object head pruning reduces accuracy marginally from 92.1% to 91.25%, and location/color head pruning has negligible effect.
These results validate the effectiveness of CCS as an interpretability tool, demonstrating that heads with high CCS scores: (i). Encode semantically aligned and task-relevant information, (ii) Are critical for model decision-making in tasks involving those concepts, and (iii). Offer targeted insights into how concepts are represented inside CLIP models. Additionally, larger models’ robustness to pruning highlights differences in representation capacity and redundancy across architectures.
5.6 Impact of CCS pruning on zero-shot video retrieval.
To further assess the importance of high CCS heads for downstream tasks, we conducted a series of zero-shot video retrieval experiments on three popular datasets: MSRVTT, MSVD, and DIDEMO under different pruning strategies. Figure  5 shows the results of this experiment. Notably, pruning high CCS (Concept Consistency Score) heads consistently leads to a substantial drop in performance across all datasets, demonstrating their critical role in preserving CLIP’s retrieval capabilities. For instance, on MSRVTT and MSVD, high CCS pruning significantly underperforms compared to low CCS and random head pruning, which show much milder performance degradation. Interestingly, low CCS and random head pruning maintain performance much closer to the original unpruned model, indicating that not all attention heads contribute equally to model competence. This consistent trend across datasets highlights that heads with high CCS scores are essential for encoding concept-aligned information necessary for accurate zero-shot video retrieval.
6 Conclusion
In this work, we proposed Concept Consistency Score (CCS), a novel interpretability metric that quantifies how consistently individual attention heads in CLIP-like models align with semantically meaningful concepts. Through extensive soft-pruning experiments, we demonstrated that heads with high CCS are essential for maintaining model performance, as their removal leads to substantial performance drops compared to pruning random or low CCS heads. Our findings further highlight that high CCS heads are not only critical for standard vision-language tasks but also play a central role in out-of-domain detection and concept-specific reasoning. Moreover, experiments on video retrieval tasks reveal that high CCS heads are crucial for capturing temporal and cross-modal relationships, underscoring their broad utility in multimodal understanding. Our study shows how CCS is a powerful interpretability metric for identifying key layers and heads in CLIP-like models.